# Multivariable Behavioral Change Modeling of Epidemics in the Presence of Undetected Infections

### v2

@@ 2.4 Estimation
with parameter vector $\Theta$ containing $\beta$, $\gamma_{1}$, $\gamma_{2}$, $\lambda$, $\phi$, $\pi^{detect}$, $k$ and $\alpha$ for the model with the multivariable alarm function. Complete data would provide $\mathbf{I}^{*}$, $\mathbf{C}^{*}$, $\mathbf{H}^{*}$, $\mathbf{R^{I}}^{*}$, $\mathbf{R^{H}}^{*}$, and $\mathbf{D}^{*}$ and initial conditions ($S_{0}$, $I_{0}$, $H_{0}$, $R_{0}$, $D_{0}$), however, from publicly available data we typically only observe $\mathbf{C}^{*}$, $\mathbf{D}^{*}$, and sometimes $\mathbf{H}^{*}$. We assume these three data sources are available and use the data-augmented MCMC approach of Lekone and Finkenstädt, (2006) to impute $\mathbf{R^{I}}^{*}$ and $\mathbf{R^{H}}^{*}$. We are not able to avoid their imputation through the IDD approach taken in the SIR model due to the multinomial transitions from the I and H compartments as the IDD specification has not yet been extended for this purpose. Data augmentation is also used to estimate the unobserved true incidence $\mathbf{I}^{*}$ in models that incorporate undetected infections. When undetected infections are ignored it is assumed that $\mathbf{I}^{*}=\mathbf{C}^{*}$ and the last component of the likelihood above is not included.
@@
All models were implemented using the R package nimble (de Valpine et al.,, 2017, 2021) with user-defined samplers for the data-augmented MCMC algorithms. For all analyses, three MCMC chains were run using various starting values of the parameters and convergence was assessed by ensuring the Gelman and Rubin diagnostic value for all parameters fell below 1.1 (Gelman et al.,, 1992). All code to reproduce the simulations and data analyses presented is available at https://github.com/ceward18/multipleDataBCM.
@@ 3.1 Simulation Set-up
A simulation study was designed to achieve the following objectives: 1) how estimation is impacted if the presence of undetected cases is ignored, 2) whether the full SIHRD model is necessary or if a simpler SIR model can still provide the same inferential results, and 3) if the multivariable alarm is estimable and how it affects model fit compared to a univariate alarm or ignoring behavioral change completely. To achieve these goals, epidemics were simulated under the full SIHRD model with undetected infections and alarm based on smoothed cases and deaths as specified in Equation 4 with smoothing over the past $m=30$ days. All epidemics were simulated in a population of one million people with five initially infected individuals and were recorded for 40 days. The detection probability was set to $\pi^{detect}=$ 25% for all simulations. Three values of the weight parameter $\alpha$ in the alarm function were used for data generation: $\alpha=0.85$ for high case importance, $\alpha=0.5$ for equal importance of cases and deaths, and $\alpha=0.15$ for high death importance. 50 epidemics were simulated in each of the three data generating scenario. All other model parameters were given the same value across the three scenarios (full specifications are provided in Web Appendix A). This led to epidemics as illustrated in Figure 1.
@@
Six models were fitted to each simulated epidemic to enumerate all possible combinations of the two compartmental structures (SIHRD and SIR) and three alarm specifications: multivariable alarm based on cases and deaths, alarm based on cases only as specified in Equation 2, and no alarm ($a_{t}=0$ for all $t$). Each of these models was fit under two assumptions: observed cases is the true incidence (ignoring undetected infections) and allowing for undetected infections to be incorporated in the modeling process as specified in Section 2.4. When undetected infections were modeled, we assume a correctly specified prior, such that $E(\pi^{detect})=0.25$ with small variance. Other priors were specified as in Section 2.4, and a full description is provided in the Web Appendix A. Three MCMC chains with different initial values were run for each model with a burn-in of 100,000 iterations and 300,000 post burn-in samples drawn every 10th iteration. Convergence was established by a Gelman and Rubin diagnostic value below 1.1 (Gelman et al.,, 1992). A small number of the SIHRD models (<< 1%) did not converge and have been excluded from the results.
@@ 3.2 Simulation Results
We use estimation of the reproductive number to assess the impact of ignoring undetected cases and to compare the SIHRD and SIR compartmental structures due to its epidemiological importance. Across simulations k=1,,Kk=1,...,K, each model’s estimated $\hat{\mathcal{R}}_{0}(t)$ was computed and compared to the true $\mathcal{R}_{0}(t)$ using the root mean squared error (RMSE) computed as $\sqrt{\frac{1}{K}\sum_{k=1}^{K}\left\{\hat{\mathcal{R}}_{0}(t)-\mathcal{R}_{0}(t)\right\}^{2}}$ at each time point $t$. In Figure 2, we present RMSE at the first and last time point to illustrate estimation prior to behavioral change occurring (start) and after behavioral change has reduced transmission (end). RMSE and the posterior distribution of $\mathcal{R}_{0}(t)$ from one simulation across all time points are provided in Web Appendix A.
@@
First, we find that models that do not incorporate any behavioral change have very poor estimation of transmission at both the start and end of the epidemics, as indicated by the large RMSE values for the models with no alarm. In assessing estimation of the reproductive number when the presence of undetected cases is ignored, we observe only minimal differences in RMSE of $\mathcal{R}_{0}(t)$, with the most improvement in RMSE seen for the SIHRD models. This result is somewhat unexpected as the true number of infections is four times higher than the observed case counts and ignoring these seems like it should negatively bias the reproductive number estimates. Inspecting the posterior distributions of all parameters in the SIHRD models (Web Appendix A Figures S4-S6), we find that while the transmission parameters ($\beta$, $k$, $\alpha$) were generally similar in distribution whether or not undetected infections were incorporated into the modeling, the parameters associated with the length of the infectious period were biased such that $\pi^{(IH)}$ was larger and $\pi^{(IR)}$ was smaller than their true values. Since hospitalizations over time are observed, ignoring undetected infections increases $\pi^{(IH)}$ as there are less infected people available to be hospitalized. In combination, the tradeoff between these two probabilities ends up having minimal impact on the estimation of $\mathcal{R}_{0}(t)$. Evaluating the posterior distributions of SIR model parameters (Web Appendix A Figures S7-S9), we see minimal changes in the posterior distributions when undetected infections were incorporated or ignored. Thus any differences in the reproductive number would be a result of differences in the number of susceptibles over time. However, as the simulated epidemics are infecting a relatively low proportion of individuals in a large population, incorporating undetected infections has only a minor impact on the total number of susceptible individuals and therefore on the estimated $\hat{\mathcal{R}}_{0}(t)$. We anticipate the incorporation of undetected infections would offer more improvement for epidemics occurring in small populations.
@@
Next we evaluate Figure 2 for differences in the RMSE of $\mathcal{R}_{0}(t)$ between the two compartmental structures and the full multivariable alarm compared to the reduced alarm based on cases only. For both modeling choices, we find that there is little penalty in using the simpler model or alarm structure in terms of estimation of $\mathcal{R}_{0}(t)$ in the high case importance and equal importance data generating scenarios. Where it becomes most important to have the correct model specification is in the scenario where deaths are highly important in driving behavior and one is interested in estimating $\mathcal{R}_{0}(t)$ towards the end of the epidemic after behavioral changes have occurred. It’s intuitive that a model ignoring deaths (i.e., the cases only model) would do poorly in this scenario where deaths are highly important. It is perhaps more interesting that even when the alarm function was correctly specified to include both cases and deaths, if the SIR compartmental structure was used $\mathcal{R}_{0}$ estimation suffers. This indicates that the more complex compartmental structure provided by the SIHRD model is important to ensure accurate estimation of the reproductive number, even when the transmission rate is correctly modeled.
@@
Finally, to evaluate the estimation of the multivariable alarm we focus on estimation of $\alpha$, the relative importance of cases in the alarm function across both the SIHRD and SIR models for all simulations (Figure 3). These results illustrate that the SIR model is biased towards cases being more important, whereas the SIHRD model does a much better job of capturing this parameter, particularly in the high deaths importance and equal importance scenarios. This is also likely the source of the increase in RMSE of $\mathcal{R}_{0}$ for the SIR model in the high deaths importance scenario observed in Figure 2. We hypothesize that this may be due to the fact that deaths are directly included into the compartmental structure in the SIHRD model. The SIHRD model does still show a small amount of bias towards higher case importance as reflected by point estimates being consistently above the true value. The estimation of $\alpha$ is worst in the equal importance scenario, but this is not necessarily surprising as we see that epidemics generated under high case importance and equal importance scenarios are more similar than epidemics generated under the high deaths importance scenario (Figure 1).
@@ 4.1 COVID-19 in Miami and Montréal
Case and death data for Miami-Dade county was obtained from the New York Times, based on reports from state and local health agencies (The New York Times,, 2024). Miami hospitalization data was obtained from the Miami-Dade COVID Project (Williams et al.,, 2021, 2024) and provides the number of COVID-19 positive patients admitted on each day. The case and death counts start on March 11, 2020 when the first case occurred in the county, however hospitalization data is not available until April 2, 2020. The missing hospitalization data was imputed using data augmentation. The population size was set at $N=$ 2,701,767, the value reported by the 2020 U.S. census. Daily case, hospitalization, and death counts for Montréal were obtained from theNational Institute of Public Health of Quebecs, (2024), with the first positive case reported on March 5, 2020. The population size for Montréal was set to $N=$ 1,762,949 as this was the value reported in the 2021 Canadian Census of Population.
@@
To illustrate the use of these models during an active outbreak, models were fit to the first six to ten weeks of each wave. The amount of data included varied by city and wave, but was chosen to include only a few weeks of data after the peak. Sensitivity analyses using different numbers of weeks of data in model fitting are presented in the Web Appendix A. As in the simulation study, six models were fitted to encompass both compartmental structures and three alarm specifications. Despite the simulation results indicating minimal benefit of incorporating undetected infections, all models incorporated this uncertainty. Prior distributions were specified as in Section 2.4 and are detailed in the Web Appendix A. Priors for the probability of detection in Miami were specified based on the Institute for Health Metrics and Evaluation (IHME) estimates of total infections and reported cases in Florida during the corresponding waves, leading to prior means of 12% in wave 1 and 34% in wave 2 (for Health Metrics and, IHME). For Montréal, these priors were based on IHME COVID-19 policy briefing on the percent of COVID-19 infections detected across Canada, with prior means of 25% detection in wave 1 and 40% detection in wave 2 (Institute for Health Metrics and Evaluation , 2022, IHME). Priors for the removal rate for hospitalized patients were chosen to reflect an average length of stay of 15 days (Alimohamadi et al.,, 2022) and priors for the removal rate and IDD curve for infected but not hospitalized individuals were chosen to reflect an average seven day infectious period before recovery. As in the simulation study, three chains were run and convergence established by a Gelman Rubin diagnostic value below 1.1.
@@ 4.2 Results
First, we evaluate model fit and compare the estimates of the relative importance of cases ($\alpha$) in informing the alarm for both cities using the SIHRD and SIR models (Table 1). We use the Widely Applicable Information Criteria (WAIC) to evaluate model fit (Watanabe and Opper,, 2010). WAIC is a likelihood-based estimate of prediction error which includes a penalty for the effective number of model parameters (Gelman et al.,, 2013). Since the SIR and SIHRD models have different likelihoods, the WAIC values are not comparable across the two compartmental structures, but it can still be used to determine the best specification of transmission rate within each structure as shown by (Ward et al.,, 2023).
During the first wave of COVID-19 cases in both cities, the estimated relative importance parameter $\alpha$ in the multivariable alarm model indicates that the majority ($\sim$88-94%) of the population alarm is due to cases during the first wave. This concurs with the WAIC values, which are essentially equivalent between the multivariable alarm and the cases only alarm in both cities. However, in Wave 2 in Miami we found that deaths were highly influential as the estimated $\hat{\alpha}$ values are close to zero and WAIC shows that model fit is significantly improved through the addition of deaths to the alarm function for both the SIHRD and SIR models. In Wave 2 in Montréal, the models estimate relatively equal contribution of case and deaths inform alarm, however the credible intervals are extremely wide. These wide intervals are likely due to the low number of deaths during the time period of model fitting (Figure 4). Again the WAIC values differ by <1<1 point between the cases and deaths alarm and the cases only alarm model. Overall the WAIC values indicate that there is little downside in terms of model fit for incorporating deaths into the alarm function even when they are not very important, and that model fit is much worse when behavioral change is not incorporated at all.
@@
Evaluating the posterior distributions of the alarm functions and reproductive numbers over time illustrates several other important findings (Figure 5). In both cities, Wave 1 corresponds with a quickly increasing alarm function, reaching an almost 70% reduction in transmission in six weeks in Miami and around 50% reduction in transmission in Montréal after eight weeks. Correspondingly, the reproductive number in both cities, which is estimated to be between 2 and 3 for both cities at the start of the pandemic decreases sharply, although it does not quite cross the threshold of one during the time period of modeling. Models which do not incorporate behavioral change through the alarm function estimate a constant reproductive number which is much lower (around 1.1 - 1.2), as these models do not contain the flexibility to allow transmission to change over time. During the second wave, the alarm functions no longer start at zero as it is not the start of the epidemic and cases and deaths are still occurring at low levels prior to the wave starting. In Miami, the alarm estimation during the second wave is heavily dependent on the inclusion of deaths in the alarm function, as our estimation of $\alpha$ showed deaths were the primary influence of behavioral change during this wave. Deaths were already elevated at the start of the wave, and alarm increases after around 40-50 days once the death rate started rapidly increasing in mid-July. The estimated reproductive number is about 1.25 in Miami and about 1.5 in Montréal at the start of Wave 2. In both cities the behavioral change models do find that $\hat{\mathcal{R}_{0}}(t)$ becomes below 1 by the end of the modeling period. As in Wave 1, the models with no alarm results in a constant reproductive number at a much lower level of transmission.

### v3

@@ 2.4 Estimation
with parameter vector $\Theta$ containing $\beta$, $\gamma_{1}$, $\gamma_{2}$, $\lambda$, $\phi$, $\pi^{detect}$, $k$ and $\alpha$ for the model with the multivariable alarm function. Complete data would provide $\mathbf{I}^{*}$, $\mathbf{C}^{*}$, $\mathbf{H}^{*}$, $\mathbf{R^{I}}^{*}$, $\mathbf{R^{H}}^{*}$, and $\mathbf{D}^{*}$ and initial conditions ($S_{0}$, $I_{0}$, $H_{0}$, $R_{0}$, $D_{0}$), however, from publicly available data we typically only observe $\mathbf{C}^{*}$, $\mathbf{D}^{*}$, and sometimes $\mathbf{H}^{*}$. We assume these three data sources are available and use the data-augmented MCMC approach of Lekone and Finkenstädtlekone to impute $\mathbf{R^{I}}^{*}$ and $\mathbf{R^{H}}^{*}$. We are not able to avoid their imputation through the IDD approach taken in the SIR model due to the multinomial transitions from the I and H compartments as the IDD specification has not yet been extended for this purpose. Data augmentation is also used to estimate the unobserved true incidence $\mathbf{I}^{*}$ in models that incorporate undetected infections. When undetected infections are ignored it is assumed that $\mathbf{I}^{*}=\mathbf{C}^{*}$ and the last component of the likelihood above is not included.
@@
All models were implemented using the R package nimble (nimble-article:2017; nimble-software:2021) with user-defined samplers for the data-augmented MCMC algorithms. For all analyses, three MCMC chains were run using various starting values of the parameters and convergence was assessed by ensuring the Gelman and Rubin diagnostic value for all parameters fell below 1.1 (gelmanrubin1992). All code to reproduce the simulations and data analyses presented is available at https://github.com/ceward18/multipleDataBCM.
2.5 Assessing Model Fit
A common approach to evaluating model fit for epidemic models is using the posterior predictive distribution. This is computed by taking a large number of draws (e.g., 10,000) from the posterior distribution of all model parameters and using each draw to simulate epidemics through forward simulation according to the specified model starting from the initial conditions. The posterior predictive distribution can be summarized over time using the mean and 95% credible intervals of the simulated epidemic trajectories. Visual comparison of the posterior predictive distribution compared to the observed data is used, with more overlap between the posterior predictive distribution and the observed data indicating better model fit. For SIR models, the observed data is the single time series of case counts, whereas for the SIHRD models, the observed data also includes time series of hospitalizations and deaths. For models with behavioral change captured in an alarm function, the alarm is dynamically updated as case and/or death counts are simulated for each posterior predicted epidemic. However, this becomes problematic for an SIR model with cases and deaths informing the alarm function as deaths are not part of the model and are subsequently not simulated as part of the posterior predictive process. Posterior prediction is only possible when the inputs to the alarm function are included in the compartmental structure, and is thus not possible for an SIR model with a multivariable alarm depending on both cases and deaths.
Another commonly used tool for model comparison is the Widely Applicable Information Criteria (WAIC) (watanabe2010asymptotic). WAIC is a likelihood-based estimate of prediction error which includes a penalty for the effective number of model parameters (gelman2013bayesian). As presented in Section 2.4, the SIR and SIHRD models have different likelihoods which means the WAIC values are not comparable across the two compartmental structures. However, within each compartmental structure, WAIC can still be used to determine the best specification of the alarm function (ward2023bayesian).
@@ 3.1 Simulation Set-up
A simulation study was designed to achieve the following objectives: 1) how estimation is impacted if the presence of undetected cases is ignored, 2) whether the full SIHRD model is necessary or if a simpler SIR model can still provide the same inferential results, and 3) if the multivariable alarm is estimable and how it affects model fit compared to a univariate alarm or ignoring behavioral change completely. To achieve these goals, epidemics were simulated under the full SIHRD model with undetected infections and alarm based on smoothed cases and deaths as specified in Equation 4 with smoothing over the past $m=30$ days. All epidemics were simulated in a population of one million people with five initially infected individuals and were recorded for 40 days. The detection probability was set to $\pi^{detect}=$ 25% for all simulations. Three values of the weight parameter $\alpha$ in the alarm function were used for data generation: $\alpha=0.85$ for high case importance, $\alpha=0.5$ for equal importance of cases and deaths, and $\alpha=0.15$ for high death importance. 50 epidemics were simulated in each of the three data generating scenarios. All other model parameters were given the same value across the three scenarios (full specifications are provided in Supplementary Table 1). This led to epidemics as illustrated in Figure 1.
@@
Six models were fitted to each simulated epidemic to enumerate all possible combinations of the two compartmental structures (SIHRD and SIR) and three alarm specifications: multivariable alarm based on cases and deaths, alarm based on cases only as specified in Equation 2, and no alarm ($a_{t}=0$ for all $t$). Each of these models was fit under two assumptions: observed cases are the true incidence (ignoring undetected infections) and allowing for undetected infections to be incorporated in the modeling process as specified in Section 2.4. When undetected infections were modeled, we assume a correctly specified prior, such that $E(\pi^{detect})=0.25$ with small variance. Other priors were specified as in Section 2.4, and a full description is provided in Supplementary Table 2. Three MCMC chains with different initial values were run for each model with a burn-in of 100,000 iterations and 300,000 post burn-in samples drawn every 10th iteration. Convergence was established by a Gelman and Rubin diagnostic value below 1.1 (gelmanrubin1992). A small number of the SIHRD models (<< 1%) did not converge and have been excluded from the results.
@@ 3.2 Simulation Results
We use estimation of the reproductive number to assess the impact of ignoring undetected cases and to compare the SIHRD and SIR compartmental structures due to its epidemiological importance. Across simulations k=1,,Kk=1,...,K, each model’s estimated $\hat{\mathcal{R}}_{0}(t)$ was computed and compared to the true $\mathcal{R}_{0}(t)$ using the root mean squared error (RMSE) computed as $\sqrt{\frac{1}{K}\sum_{k=1}^{K}\left\{\hat{\mathcal{R}}_{0}(t)-\mathcal{R}_{0}(t)\right\}^{2}}$ at each time point $t$. In Figure 2, we present RMSE at the first and last time point to illustrate estimation prior to behavioral change occurring (start) and after behavioral change has reduced transmission (end). RMSE and the posterior distribution of $\mathcal{R}_{0}(t)$ from one simulation across all time points are provided in Supplementary Figures 1 and 2.
@@
First, we find that models that do not incorporate any behavioral change have very poor estimation of transmission at both the start and end of the epidemics, as indicated by the large RMSE values for the models with no alarm. In assessing estimation of the reproductive number when the presence of undetected cases is ignored, we observe only minimal differences in RMSE of $\mathcal{R}_{0}(t)$. This is consistent with previous literature indicating that the estimation of the reproductive number is unaffected by under-detection when there is a constant detection probability (cori2013new; gostic2020practical). Inspecting the posterior distributions of all parameters in the SIHRD models (Supplementary Figures 4-6), we find that while the transmission parameters ($\beta$, $k$, $\alpha$) were generally similar in distribution whether or not undetected infections were incorporated into the modeling, the parameters associated with the length of the infectious period were biased such that $\pi^{(IH)}$ was larger and $\pi^{(IR)}$ was smaller than their true values. Since hospitalizations over time are observed, ignoring undetected infections increases $\pi^{(IH)}$ as there are less infected people available to be hospitalized. In combination, the tradeoff between these two probabilities ends up having minimal impact on the estimation of $\mathcal{R}_{0}(t)$. Evaluating the posterior distributions of SIR model parameters (Supplementary Figures 7-9), we see minimal changes in the posterior distributions when undetected infections were incorporated or ignored. Thus any differences in the reproductive number would be a result of differences in the number of susceptibles over time. However, as the simulated epidemics are infecting a relatively low proportion of individuals in a large population, incorporating undetected infections has only a minor impact on the total number of susceptible individuals and therefore on the estimated $\hat{\mathcal{R}}_{0}(t)$.
@@
We also evaluate Figure 2 for differences in the RMSE of $\mathcal{R}_{0}(t)$ between the two compartmental structures and the full multivariable alarm compared to the reduced alarm based on cases only. For both modeling choices, we find that there is little penalty in using the simpler model or alarm structure in terms of estimation of $\mathcal{R}_{0}(t)$ in the high case importance and equal importance data generating scenarios. Where it becomes most important to have the correct model specification is in the scenario where deaths are highly important in driving behavior and one is interested in estimating $\mathcal{R}_{0}(t)$ towards the end of the epidemic after behavioral changes have occurred. It’s intuitive that a model ignoring deaths (i.e., the cases only model) would do poorly in this scenario where deaths are highly important. It is perhaps more interesting that even when the alarm function was correctly specified to include both cases and deaths, if the SIR compartmental structure was used $\mathcal{R}_{0}$ estimation suffers. This indicates that the more complex compartmental structure provided by the SIHRD model is important to ensure accurate estimation of the reproductive number, even when the transmission rate is correctly modeled.
Next, we evaluate model fit using the posterior predictive distributions as described in Section 2.5. The posterior predictive distributions for one randomly selected simulated epidemic from each data generating scenarios are presented in Figure 3. As we would expect given the poor estimation of $\mathcal{R}_{0}(t)$ for models that do not incorporate behavioral change, we find that these models also fit very poorly. Transmission is severely underestimated leading to posterior predictive distributions where on average less than 10 individuals are infected across both the SIHRD and SIR compartmental structures and all three data generating scenarios. Also as expected, the posterior predictive distributions for the SIHRD model with the full multivariable alarm are fairly close to the truth across all three data generating scenarios. When the alarm is specified based on cases only, the SIHRD model fits well in scenarios of high case importance or equal importance but significantly underestimates the peak and overestimates the tail when deaths are highly important. Similar trends are seen for the SIR model with cases only in the alarm function, and this model also seems to overestimate the peak in the high case importance scenario.
Figure 3: Posterior predictive means (dashed lines) and 95% credible intervals (shaded regions) for one randomly selected simulation compared to true cases (solid black), hospitalizations (solid gold), and deaths (solid blue).
@@
Finally, to evaluate the estimation of the multivariable alarm we focus on estimation of $\alpha$, the relative importance of cases in the alarm function across both the SIHRD and SIR models for all simulations (Figure 4). These results illustrate that the SIR model is biased towards cases being more important, whereas the SIHRD model does a much better job of capturing this parameter, particularly in the high deaths importance and equal importance scenarios. This is also likely the source of the increase in RMSE of $\mathcal{R}_{0}$ for the SIR model in the high deaths importance scenario observed in Figure 2. We hypothesize that this may be due to the fact that deaths are directly included into the compartmental structure in the SIHRD model. The SIHRD model does still show a small amount of bias towards higher case importance as reflected by point estimates being consistently above the true value. The estimation of $\alpha$ is worst in the equal importance scenario, but this is not necessarily surprising as we see that epidemics generated under high case importance and equal importance scenarios are more similar than epidemics generated under the high deaths importance scenario (Figure 1) which may indicate that it is harder to differentiate between moderate and high values of $\alpha$ than between moderate and low values of $\alpha$.
@@ 4.1 COVID-19 in Miami and Montréal
Case and death data for Miami-Dade county was obtained from the New York Times, based on reports from state and local health agencies (NYtimesdata). Miami hospitalization data was obtained from the Miami-Dade COVID Project (williams2021lessons; miamicovidproject) and provides the number of COVID-19 positive patients admitted on each day. The case and death counts start on March 11, 2020 when the first case occurred in the county, however hospitalization data is not available until April 2, 2020. The missing hospitalization data was imputed using data augmentation. The population size was set at $N=$ 2,701,767, the value reported by the 2020 U.S. census. Daily case, hospitalization, and death counts for Montréal were obtained from the National Institute of Public Health of Québec montrealData, with the first positive case reported on March 5, 2020. The population size for Montréal was set to $N=$ 1,762,949 as this was the value reported in the 2021 Canadian Census of Population.
@@
To illustrate the use of these models during an active outbreak, models were fit to the first six to ten weeks of each wave. The amount of data included varied by city and wave, but was chosen to include only a few weeks of data after the peak. Sensitivity analyses using different numbers of weeks of data in model fitting are presented in Supplementary Table 5. As in the simulation study, six models were fitted to encompass both compartmental structures and three alarm specifications. Despite the simulation results indicating minimal benefit of incorporating undetected infections, all models incorporated this uncertainty. Prior distributions were specified as in Section 2.4 and are detailed in Supplementary Tables 3 and 4. Priors for the probability of detection in Miami were specified based on the Institute for Health Metrics and Evaluation (IHME) estimates of total infections and reported cases in Florida during the corresponding waves, leading to prior means of 12% in wave 1 and 34% in wave 2 (ihme_miami). For Montréal, these priors were based on IHME COVID-19 policy briefing on the percent of COVID-19 infections detected across Canada, with prior means of 25% detection in wave 1 and 40% detection in wave 2 (ihme_canada). Priors for the removal rate for hospitalized patients were chosen to reflect an average length of stay of 15 days (alimohamadi2022hospital) and priors for the removal rate and IDD curve for infected but not hospitalized individuals were chosen to reflect an average seven day infectious period before recovery. As in the simulation study, three chains were run and convergence established by a Gelman Rubin diagnostic value below 1.1.
@@ 4.2 Results
First, we evaluate model fit using WAIC and compare the estimates of the relative importance of cases ($\alpha$) in informing the alarm for both cities using the SIHRD and SIR models (Table 1). We find that, in general, estimates of $\alpha$ are similar between the two approaches, but the SIHRD model tends to attribute more behavioral change to deaths than the SIR model, as we would expect based on the results of our simulation study. During the first wave of COVID-19 cases in both cities, the estimated relative importance parameter $\alpha$ in the multivariable alarm model indicates that the majority ($\sim$88-94%) of the population alarm is due to cases during the first wave. This concurs with the WAIC values, which are essentially equivalent between the multivariable alarm and the cases only alarm in both cities. However, in Wave 2 in Miami we found that deaths were highly influential as the estimated $\hat{\alpha}$ values are close to zero and WAIC shows that model fit is significantly improved through the addition of deaths to the alarm function for both the SIHRD and SIR models. In Wave 2 in Montréal, the models estimate relatively equal contribution of case and deaths inform alarm, however the credible intervals are extremely wide. These wide intervals are likely due to the low number of deaths during the time period of model fitting (Figure 5). Again the WAIC values differ by <1<1 point between the cases and deaths alarm and the cases only alarm model. Overall the WAIC values indicate that there is little downside in terms of model fit for incorporating deaths into the alarm function even when they are not very important, and that model fit is much worse when behavioral change is not incorporated at all.
@@
As WAIC can only provide insight into relative specification of the alarm function within a compartmental structure, we also use the posterior predictive distribution of cases (and hospitalizations and deaths for the SIHRD models) trajectories as a global comparison of model fit across all models. (Figure 6). As expected, the posterior predictive distributions of the models with no alarm do not resemble the observed data as transmission is severely underestimated. Comparing the SIHRD models to the SIR models, we find the SIHRD models provide better fit to Wave 1 in both cities, although the posterior prediction of cases for the SIHRD models still tends to underestimate the initial increase in cases and estimate a slightly delayed peak. In Wave 2, the SIR cases only model has the most overlap between the posterior predictive distribution and the observed case counts. This is not surprising as our simulation study also found this model to have adequate posterior predictive fit until later into the wave even when it was not the true data generating model. Of note, the posterior predictive distribution of cases in Wave 2 in Miami for the SIHRD model with cases and deaths in the alarm function does not match the observed data, despite having a significantly lower WAIC value then the SIHRD model with only cases informing the alarm. Additionally for this wave, the SIR cases only model has a posterior predictive distribution that best matches the observed data. Although the posterior predictive distribution of cases was poor, both SIHRD models had reasonable posterior predictive coverage of the observed hospitalization and death curves (see Supplementary Figure 10 for the zoomed in posterior predictive distributions of hospitalizations and deaths). Combined, these results may indicate that the strong prior on the constant hospitalization rate, which is needed for incorporating the undetected infections is misspecified. Further discussion on this assumption and the feasibility of potential alternate specifications is provided in Section 5.
Figure 6: Posterior predictive means (dashed lines) and 95% credible intervals (shaded regions) for cases, hospitalizations, and deaths compared to observed cases (solid black), hospitalizations (solid gold), and deaths (solid blue) for two COVID-19 waves in Miami and Montréal. Hospitalizations and deaths are only shown for the SIHRD model as posterior prediction of deaths is not possible from the SIR model (see Section 2.5). Hospitalization data for Wave 1 in Miami was only available from April 2, $20$ and was otherwise imputed.
@@
Lastly, we evaluate the posterior distributions of the alarm functions and reproductive numbers over time, which illustrates several other important findings (Figure 7). In both cities, Wave 1 corresponds with a quickly increasing alarm function, reaching an almost 70% reduction in transmission in six weeks in Miami and around 50% reduction in transmission in Montréal after eight weeks. Correspondingly, the reproductive number in both cities, which is estimated to be between 2 and 3 for both cities at the start of the pandemic decreases sharply, although it does not quite cross the threshold of one during the time period of modeling. Models which do not incorporate behavioral change through the alarm function estimate a constant reproductive number which is much lower (around 1.1 - 1.2), as these models do not contain the flexibility to allow transmission to change over time. During the second wave, the alarm functions no longer start at zero as it is not the start of the epidemic and cases and deaths are still occurring at low levels prior to the wave starting. In Miami, the alarm estimation during the second wave is heavily dependent on the inclusion of deaths in the alarm function, as our estimation of $\alpha$ showed deaths were the primary influence of behavioral change during this wave. Deaths were already elevated at the start of the wave, and alarm increases after around 40-50 days once the death rate started rapidly increasing in mid-July. The estimated reproductive number is about 1.25 in Miami and about 1.5 in Montréal at the start of Wave 2. In both cities the behavioral change models do find that $\hat{\mathcal{R}_{0}}(t)$ becomes below 1 by the end of the modeling period. As in Wave 1, the models with no alarm results in a constant reproductive number at a much lower level of transmission.